OpenStudy (anonymous):

At 350 degrees C, nitrogen has a velocity of 800 m/s. Find the velocity of hydrogen at the same temperature. I understand that the equation for Graham's law is R1/R2 = square root of M2/M1 I just can't seem to figure out how to set up this equation.

OpenStudy (anonymous):

don't think this is graham's law. That law has to do with effusion, and here you're most likely talking about root mean square

OpenStudy (anonymous):

\[v = \sqrt{\frac{ 3RT }{ M }}\]

OpenStudy (anonymous):

The above is the equation that you have to use. First use the nitrogen mass to solve for the temperature. You have everything to do this. v is the velocity 800m/s. R is the constant. M is the molar mass (be careful you have to use N2 here).

OpenStudy (anonymous):

after you're done, plug that same temperature, and use the mass of H2 to find the velocity for hydrogen

OpenStudy (aaronq):

Graham's law ends up as the same thing: \(\sf \large v=\sqrt{\dfrac{3RT}{M}}\rightarrow \dfrac{v}{\sqrt{\dfrac{3RT}{M}}}=0\) \(\sf \large \dfrac{v_1}{\sqrt{\dfrac{\cancel{3RT}}{M_1}}}=\dfrac{v_2}{\sqrt{\dfrac{\cancel{3RT}}{M_2}}} \rightarrow v_1\sqrt M_1= v_2\sqrt M_2\) \(\sf \large \dfrac{v_1}{v_2}=\dfrac{\sqrt M_2}{\sqrt M_1}=\sqrt{\dfrac{M_2}{M_1}}\)

OpenStudy (anonymous):

Huh, never thought about it that way. I only vaguely remember the formula in high school in AP Chem and we didn't make any connections like this